Find three consecutive integers such that the sum of the largest and 5 times the smallest is -244?

Answer by mananth(16946)

let integers be x , x+1, x+2
x+2+5x=-244
6x+2=-244
6x=-246
x=-246/6
x=-41
the numbers are -41,-40,-39